Gillette's most recent annual dividend was $8 per share. The company expects the growth of its dividends to be stable at 3% per
year going forward.
a) If investors require a 9% return, what is the current value of Gillette's stock? (round to nearest cent)
b) If the stock currently trades at $116.57 per share, what is the dividend growth rate investors expect? (round to nearest percent)
      Hint: When the constant-growth formula is solved for the growth variable, it is g = -D/P + r
Answer #1

a. Price = Div1/(r - g)

Div1 = 8 * (1 + 0.03)

Div1 = 8.24

Price = 8.24/(0.09 - 0.03)

Price = $137.3333333333

b. Price = Div1/(r - g)

r - g = Div1/Price

g = r - Div1/Price

g = 0.09 - 8.24/116.57

g = 0.01931285923

g ~ 2%

Can you please upvote? Thank You :-)